A patient with a herpes simplex infection is to receive Acyclovir 325mg. The medication is provided by the pharmacy as 325mg/50mL 0.9% Sodium Chloride. The medication should be infused over 1 hour Q8H using a volumetric infusion pump. The tubing drop factor is 20 gtt/mL.

The nurse should set the pump at what rate in mL per hour?
Answer #1

ANSWER.

* Calculation of rate in ml/hr.

       Rate(ml/hr) = Volume(ml) /Time(hr).

     Calculation of volume.

           Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

        Desired dose(D) = 325 mg.

        Dose in hand (H) = 325 mg.

             Quantity (Q) = 50 ml.

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                       = 325/325 * 50

                                                       = 1 * 50

                                                       = 50 ml.

                     Time(hr) = 1 hr.

         So

               Rate(ml/hr) = 50 ml/hr.